OpenStudy (anonymous):

What is the solution to the rational equation 4/3-2/x-5=6/3x-15

OpenStudy (anonymous):

a) x = -8 b) x = -2 c) x=2 d) x=8

hero (hero):

\[\frac{4}{3} - \frac{2}{(x - 5)} = \frac{6}{3x - 15}\] Factor 3x - 15 to get: \[\frac{4}{3} - \frac{2}{(x - 5)} = \frac{6}{3(x - 5)}\] 6\3 = 2 so \[\frac{4}{3} - \frac{2}{(x - 5)} = \frac{2}{(x - 5)}\] Add 2/(x - 5) to both sides: \[\frac{4}{3} = \frac{2}{(x - 5)} + \frac{2}{(x - 5)}\] \[\frac{4}{3} = \frac{4}{(x - 5)} \] Divide both sides by 4; Multiply both sides by 3 to get: \[\frac{4}{4} = \frac{3}{(x - 5)} \] \[1 = \frac{3}{(x - 5)} \] Multiply both sides by x - 5 to get \[x - 5 = 3 \] Finish it.

OpenStudy (anonymous):

x=8